### v1

# Tianhao Wei∗,1, Liqian Ma∗,1,2, Rui Chen1, Weiye Zhao1, Changliu Liu1

Abstract
The requirements for real-world manipulation tasks are diverse and often conflicting; some tasks necessitate force constraints or collision avoidance, while others demand high-frequency feedback. Satisfying these varied requirements with a fixed state-action representation and control strategy is challenging, impeding the development of a universal robotic foundation model. In this work, we propose Meta-Control, the first LLM-enabled automatic control synthesis approach that creates customized state representations and control strategies tailored to specific tasks. Meta-Control leverages a generic hierarchical control framework to address a wide range of heterogeneous tasks. Our core insight is the decomposition of the state space into an abstract task space and a concrete tracking space. By harnessing LLM’s extensive common sense and control knowledge, we enable the LLM to design these spaces, including states, dynamic models, and controllers, using pre-defined but abstract templates. Meta-Control stands out for its fully model-based nature, allowing for rigorous analysis, efficient parameter tuning, and reliable execution. It not only utilizes decades of control expertise encapsulated within LLMs to facilitate heterogeneous control but also ensures formal guarantees such as safety and stability. Our method is validated both in real-world scenarios and simulations across diverse tasks with conflicting requirements, such as collision avoidance versus convergence and compliance versus high precision. Videos and additional results are at meta-control-paper.github.io
Keywords: Embodied agent, Model-based Control, LLM, Manipulation
Figure 1: Open world robot tasks have inherently different and even opposite requirements that can be difficult to satisfy holistically with fixed state-action representations and fixed control strategies. To address this problem, Meta-Control leverages hierarchical formulation and LLMs, enabling automatic control synthesis by automatically customizing the most appropriate state action representations, dynamic models, and controllers.
1 Introduction
Recent advances in robot learning have enabled embodied agents to interact intelligently with the environments via methods such as predefined action primitives (i.e. skills) [1, 2, 3], parameterized policies [4, 5, 6], and reinforcement learning (RL) [7, 8, 9].
However, existing methods face a common challenge: real-world tasks exhibit inherently diverse and often conflicting requirements that are difficult to satisfy with a fixed control strategy. For instance, as illustrated in fig. 1, a pick-and-place task requires precise movements to ensure collision avoidance, which requires safe position control. Conversely, position control is unsuitable for an open-door task due to the difficulty of planning a trajectory that perfectly aligns with the door’s swing path, making compliant control more favorable. Similarly, pole balancing tasks demand controllers that ensure the pole’s convergence (a position-attracting goal) that opposes collision avoidance (a position-avoiding goal). On the other hand, although visuomotor methods can potentially address heterogeneous tasks, they often lack reliability and explainability. Models with predefined action pools may cover different task types but require manual construction and have limited applicability, hindering their scalability to diverse tasks. These challenges restrict existing methods from generalizing to various open-world manipulation tasks with varying constraints.
To overcome these limitations, we propose Meta-Control, a novel method that automatically customizes robot skills for diverse challenges of open world tasks. Our core insights are: 1) While it is difficult to overcome arbitrary challenges with a fixed representation and control strategy, each specific challenge can be effectively tackled with a customized representation and control strategy; 2) Customization done by LLM can benefit from centuries of practice in control system engineering. For example, the agent would select a compliant controller for opening a door, and a hybrid position/force controller for board wiping tasks.
Motivated by these observations, we formulate robot skill design as a control system synthesis problem and leverage LLMs to design the most appropriate representation and control strategy for a given task. We inherit the hierarchical design scheme and standardize it to fit the control synthesis framework. The hierarchy involves a high-level task space and a low-level tracking space. The task space is an abstracted, intuitive space for accomplishing the task (e.g., Cartesian space or gripper pose space) while the tracking space usually represents the robot state space where low-level constraints can be specified and task-level commands are followed. The LLM determines both the task and tracking spaces, along with the corresponding dynamic models and controllers for each space. In general, the task controller focuses on high-level objectives, while the tracking controller emphasizes low-level control with constraint satisfaction. This hierarchical design significantly enhances the capability of synthesized control systems compared to motion primitive methods.
Meta-Control offers several benefits: 1. It enables the synthesis of challenging heterogeneous robotic skills for unseen tasks, allowing each task to be accomplished with the most suitable representation and control system tailored to task-specific requirements. 2. Unlike previous work that primarily utilizes spatial priors from LLMs (e.g., object localization), Meta-Control leverages the internalized control knowledge of LLMs which encompasses decades of modeling and optimization efforts in the form of control system design. 3. The synthesized control system is fully model-based which allows rigorous analysis, efficient parameter tuning, and formal guarantees (e.g., safety and stability), leading to reliable and trustworthy execution compared to end-to-end methods.
2 Related work
Recent advancements in designing embodied agents for diverse tasks can be broadly classified into four categories: skill libraries, parameterized policies, end-to-end methods, and representation learning. Meta-Control integrates aspects of these approaches. A Meta-Control synthesized control system can be viewed as an end-to-end parameterized policy composed hierarchically of unit primitives. The representation is chosen by LLM. The parameters are inferred by LLM from trial and error, akin to RL with an LLM provided reward. By integrating the strengths of these approaches, Meta-Control sets a new standard for synthesizing robot skills, offering a flexible, robust, and explainable solution for a wide range of real-world tasks.
Skill library allows diverse control strategies and composite execution through predefined libraries of motion primitives with high-level APIs. These libraries can be dynamically called or combined by the LLM to accomplish tasks [1, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] Other works build skill libraries through behavior cloning, reinforcement learning, or bootstrapping [5, 20, 10, 21, 22, 23, 24, 25, 26, 19, 27, 28, 29]. However, these methods require manual construction and are often limited to specific task types (e.g., pick and place). In contrast, Meta-Control can synthesize new, heterogeneous skills on the fly using predefined dynamic models and controller templates.
Parameterized policy differs from the skill library methods in using a fixed strategy, such as MPC or RL with varying parameters to compose policies for different tasks. The parameters may include dynamic models [30, 31, 32, 33, 34, 35], constraints [36, 4, 37], or costs / objectives / rewards [38, 39, 40, 41, 7, 8]. These parameters can be inferred by LLMs [4, 41, 7, 8], or learned from data [30, 38, 39]. Another typical case is a hierarchical policy in which the parameters are high-level commands for an instruction following controller [42, 29]. A notable limitation is that the chosen parameterized policy inherently constrains the method’s capability. For example, an MPC that generates end-effector actions in Cartesian space cannot produce force-compliant control in joint space. RL-based methods that learn parameters from rewards are unsuitable for online skill synthesis. In comparison, Meta-Control can dynamically generate parameterized policies and is highly flexible.
End-to-End models directly map perceptions to robot actions. [6, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] leverage visual language models (VLMs) to translate instructions and observations to robot arm actions. Besides explicitly mapping observation and instruction to action, it’s also admissible to first learn energy function [53], pixel-wise or voxel-wise affordance [54, 55, 56], and then extract actions from that energy function or affordance map. Recent works utilize diffusion models to learn from demonstration [57, 58, 20], enabling multimodal action distribution. End-to-end methods often output Cartesian actions (position, orientation, velocity, etc.), requiring additional motion planning to generate joint-level movement, and therefore inherently limit the capability. Furthermore, the training data domain limits the generalizability. Meta-Control overcomes these limitations by integrating LLM with model-based control strategies, ensuring robustness and explainability.
Representation learning maps high-dimensional observations into low-dimensional representations that are more semantically meaningful and easier to control with [59, 60, 61]. It can be combined with various control strategies [62]. Efforts to develop generic representations for different tasks include task-specific representation [63, 64], visual representation [65, 61, 66, 67, 68, 69, 70, 71], and multi-modal representations [72, 73, 55, 67, 74, 75]. However, learned representations often require fixed raw inputs (e.g., camera input) and lack explainability. In contrast, Meta-Control selects a subset of available observations for representation, enhancing explainability and adaptability to different inputs.
3 Method
Figure 2: Overview of Meta-Control. The skill description is given by the user. Other prompts are automatically generated based on the system setup and code base. The LLM designs the system with dynamic model and controller templates, connects all modules by aligning the interfaces, then evaluates the synthesized control system and optimizes parameters based on execution results and optional human feedback.
In this work, we focus on synthesizing robot skills, defined as unit actions (e.g., grab the eraser, erase the marks) from robot tasks instructed via free-form language $\mathcal{L}$ (e.g., clean the whiteboard) [4]. We assume that the decomposition from task to skills: $\mathcal{L}\to{\ell_{1},\ell_{2},\dots,\ell_{n}}$ is given by a task-level planner, which can be LLM-based or search-based. Our focus is to synthesize a control system to accomplish each of the skills described by $\ell_{i}$.
Control system synthesis is a fundamental aspect of control engineering that involves designing and implementing control systems to manage the behavior of dynamic systems. The goal of control system synthesis is to ensure that a system operates as desired, achieving specified performance criteria such as stability, accuracy, and efficiency. The process typically involves two key steps: system modeling and controller synthesis. Directly designing the control system for open-world skills is very difficult because the design space is infinite. Existing work usually simplifies the process in different ways as discussed in section 2. To address these challenges, we present a hierarchical control formulation for skill synthesis and then introduce the three key procedures for control system synthesis with LLM: design via templates, interface alignment, and parameter optimization.
3.1 Generic Hierarchical Control Framework
Our key insight is that a bilevel hierarchical control framework can represent various skills. Given a skill description $\ell_{i}$, we can synthesize a high-level task controller in an intuitive abstract space and a low-level tracking controller that tracks the high-level control in the robot state space. We formulate the problem in three spaces: state space, measurement space, and task space.
[…]
where $x$ is the system state, $u$ is the state space control input, $y$ is the output or measurement of the system, $z$ is the state of the task space, and $v$ is the task space control input. The task space contains intuitive and high-level states, such as the gripper poses for robot arms and the center-of-mass for quadrupeds. $y$ has to be measurable or extractable from perception. $f,g,h$ are dynamic models of the corresponding system. The task space can sometimes be omitted when it coincides with the state space, such as when a robot arm skill directly specifies joint goals. We denote the task space controller by $\pi_{v}(y)$, and the tracking space controller by $\pi_{u}(y,v)$. Essentially, we decompose skill synthesis into two subproblems with $J$ denoting objectives and $c$ denoting constraints:
$\begin{matrix}&\min_{\pi(y)}J(x(t),u(t))&\implies&\min_{\pi_{v}(y)}J_{z}(z(t),v(t))&\&&\min_{\pi_{u}(y,v)}J_{x}(x(t),u(t))\\ &s.t.~{}c(x)\leq 0,&&s.t.~{}c_{z}(z)\leq 0,&&s.t.~{}c_{x}(x)\leq 0\end{matrix}$
The first sub-problem solves the task control input $v$ under task-space constraint $c_{z}$. The second sub-problem tracks $v$ by solving the state control input $u$ under state-space constraint $c_{x}$. Our goal is to use LLM to 1) implicitly design the proper task state $z$; 2) explicitly design the dynamic models $h(z,v)$ and $f(x,u)$; 3) implicitly design the objective $J_{z}$ and $J_{x}$ and the constraints $c_{z}(z)$ and $c_{x}(x)$. 4) explicitly design the task and tracking controller $\pi_{v}(y)$, $\pi_{u}(y,v)$. 5) Adjust the parameters for $h$, $f$, $\pi_{v}(y)$, and $\pi_{u}(y,v)$ to achieve the desired performance.
Figure 3: Comparison of Meta-control and trajectory planning based method on real robot for wiping board and opening door. The trajectory based method fails to erase the mark because it neglects force requirement. Furthermore, opening a door with a trajectory-based method leads to displacement of the cabinet because the planned trajectory does not precisely align with the door’s movement. That may damage the door if the cabinet is fixed. In contrast, Meta-Control addressed these challenges with properly customized control systems.
3.2 Control System Synthesis by LLM
Designing $z$, $h$, $f$, $\pi_{v}$, and $\pi_{u}$ directly is challenging due to the infinite possible spaces and dynamic models, the need for accurate and deep understanding of the robotic system and data flow, and the dependencies between different modules. To address these challenges, we define model and controller templates to constrain the form of dynamic models and controllers. Then, an LLM synthesizes the control system in three stages: design via templates, interface alignment, and parameter optimization (see fig. 2).
Design via Templates The LLM designs the control system by selecting the appropriate models and controllers from the template library provided in the prompt. Templates are predefined object classes that need to be instantiated with concrete arguments. For example, we offer a dynamic model template called LinearModel which can be instantiated by passing in four matrices $A,B,C,D$ describing a linear system, and a controller template LQRController requires matrices $Q,R$ and vectors $x_{0},u_{0}$. The LLM chooses whether a template is for the task space or the tracking space. Templates include functionality descriptions, argument requirements, and interface format. The design also relies on available measurements and expected control signals, which vary across tasks.
Interface Alignment The LLM connects the instantiated modules to each other and the robot by matching the outputs and inputs. Specifically, the LLM has to implement 1) a task input converter to compose the task controller input based on available measurements $y$; 2) a tracking input converter to compose the tracking controller input from the task controller’s output $v$ and available measurements $y$. For example, in the balance cart pole skill, the task controller LQRController takes a 4D vector as input, requiring the task input converter to extract these states from more than 20 available measurements. Then, the LQRController computes the desired force on the cart, which is a scalar, while the downstream tracking HybridPositionForceController requires two 6D vectors as input for desired gripper position and force. The tracking input converter must prepare the input by padding the task controller output and combining with available measurements.
Parameter Optimization The LLM infers unknown parameters for the dynamic models and controllers via trial and error. The LLM identifies which parameters need tuning, selects performance metrics from measurements, and adjust the parameters over multiple runs to improve performance. For example, in the balance cart-pole skill, LLM identifies $Q$ and $R$ are parameters and cart position and pole angle are metrics. This process is efficient due to the well-studied parameter-performance relationships in common dynamics and controllers, which are internalized by LLMs.
4 Experiment
Figure 4: Three manipulation tasks that have inherently different challenges and requirements. The balance task requires an accurate, high-frequency, feedback controller. The open door task requires properly handling articulated objects, the executed trajectory has to perfectly match the swing path. The safe pick and place task requires guaranteeing collision avoidance for the whole robot arm.
Figure 5: Meta-Control can automatically identify hyper-parameters that require tuning and tune them to accomplish challenging tasks. The figure shows the trajectory of the arm-held cart-pole system before and after tuning the synthesized controller. The hyper-parameters $Q=$diag$(10,1,100,1)$, $R=0.01$ are chosen and tuned by the LLM with only 2 rounds of trial-and-error.
The experiment is designed to manifest the following features: 1) Meta-Control enables synthesizing skills for heterogeneous requirements; 2) Meta-Control exploits control knowledge from LLM; 3) The Meta-Control pipeline is critical to the success; 4) Meta-Control works on a wide range of open world tasks; 5) Meta-Control enables model-based analysis, providing formal guarantees on various properties. 6) Meta-Control transfers to real robots and different embodiments easily.
We implement our pipeline with Drake[76], a framework designed for model-based control. For the hardware experiment, we used a Kinova Gen3 robot arm. The language model is GPT 4.0.
4.1 Meta control enables synthesizing challenging skills
As shown in fig. 3 and fig. 4, Meta-Control successfully synthesizes controllers for various challenging tasks with inherently different requirements, both in simulation and the real world. The task challenges and synthesized control systems are summarized in table 1. A more detailed description can be found in appendix B. fig. 5 shows that Meta-Control can identify core metrics and critical parameters that affect performance, then efficiently and effectively tune the parameters based on the metric through trial and error. The complete process of skill synthesis can be found at appendix C.
Task Challenge Meta-Control designed system
Open the door The robot trajectory must perfectly align with the door’s swing path. position control can easily lead to damage or failure. CartesianTrajectoryController + CartesianStiffnessController, allowing imperfect trajectory planning and tracking with compliant behavior to avoid damage or failure.
Balance the cart pole The pole is non-actuated. The system is sensitive, requiring high-frequency feedback and convergence guarantee. LQRController + HybridPositionForceController. The LQR controller gives the force to be applied on the cart, and the hybrid position/force controller tracks the desired force on the y-axis while maintaining a neutral pose on the x-axis and the z-axis.
Collision-free pick and place whole-body collision free in continuous time during the whole task. KinematicTrajectoryMPC + SafeController, allowing discrete-time planning and continuous-time whole-body collision-free tracking.
Wipe the whiteboard Two different objectives: tracking position and maintaining force CartesianTrajectoryController + HybridPositionForceController, allowing position tracking while maintaining a desired force on the whiteboard.
Table 1: Experiment tasks, challenges and Meta-Control synthesized controllers..
Figure 6: Meta-Control generalizes to different embodiments because the synthesized controller is fully model-based. A controller synthesized on Kinova can transfer to Franka Pranda simply by replacing the robot dynamic model.
Figure 7: Meta-Control synthesized control systems are robust to attribute/state changes because of the model-based design.
Method API API + Formulation API + Template Meta-Control
Balance design 30% 90% 60% 100%
implementation 0% 30% 20% 90%
execution 0% 0% 0% 70%
Open door design 40% 50% 60% 100%
implementation 10% 20% 10% 100%
execution 0% 0% 0% 80%
Safe Pick&place design 0% 0% 40% 90%
implementation 0% 0% 0% 90%
execution 0% 0% 0% 90%
Table 2: The ablation study shows that hierarchical formulation (Formulation) and simplification by dynamic model and controller templates (Templates) both improves the success rate, compared to directly synthesis with only the low-level API.
4.2 Meta control exploits dynamics priors
In contrast to previous work that focuses more on spatial relationship priors [1, 4, 5], Meta-Control exploits the LLM’s internalized knowledge of dynamics, enabling to synthesize high-performance controllers rigorously. In the balance cart pole task, LLM designs the task space dynamics $h(z,v)$ with a linear approximation in the form of $\dot{z}=Az+Bv$ around the upright position of the pole, where $A$ and $B$ are directly given by the LLM: $A=[[0,1,0,0];[0,0,\frac{m_{\text{pole}}g}{m_{\text{cart}}},0];[0,0,0,1];[0,0,\frac{g(m_{\text{cart}}+m_{\text{pole}})}{l_{\text{pole}}m_{\text{cart}}},0]]$, $B=[0;\frac{1}{m_{\text{cart}}};0;-\frac{1}{l_{\text{pole}}m_{\text{cart}}}]$, where $m$ denotes mass, $l$ denotes length, $g$ is the acceleration of gravity. Detailed description of the synthesized control system can be found at appendix B.
4.3 Ablation study of the Meta Control pipeline
To demonstrate the necessity of the hierarchy formulation and simplification using templates, we show the success rate of control system synthesis under different conditions. We test the success rate in 3 stages: design (via templates), implementation (interface alignment), and execution. We say that a design is successful if the LLM-designed control system has the potential to perform the skill judged by an expert. We say that the implementation is successful if the system can run without errors. We say that the execution is successful if the control system finishes the skill as desired. We repeat each task $10$ times to compute the success rate. Randomness is caused by the LLM with a default temperature $1.0$. As shown in table 2, we can see that the success rates of all steps in the baseline are lower than Meta-Control. Although LLM gives reasonable architectures to finish the task in the Balance and Open Door task, the LLM fails to provide correct code and parameters to realize the control system due to the complexity of the system and the huge design space. With all the modules, we achieve the highest success rate for all tasks. The Safe Pick&Place task is especially difficult for the baseline because the baseline methods, even though prompted to avoid collision, were unable to successfully design a controller that can avoid collision continuously.
4.4 Generalization to different attributes/states
Meta-Control synthesized control systems can easily generalize to scenarios of different attributes/states due to the model-based nature. Given a successfully synthesized control system, we test different attributes/states and calculate the success rate. The range of change and the results are shown in table 3. Examples are shown in fig. 7. Meta-Control achieved a 100% success rate for all scenarios.
4.5 Meta Control enables Model-based Analysis
One merit of our method is that the synthesized controller is fully model-based, allowing a rigorous formal analysis for a variety of properties:
Convergence and Stability: For the balance cart-pole task, we can provide a guarantee of convergence by solving the Riccati equation for the LQR controller. The closed-loop system matrix $A-BK$ has the following four eigenvalues: $-412.29,-9.925,-1.502+1.175j,-1.502-1.175j.$ All of them have negative real parts, which means that the system is guaranteed to converge. More rigorous analysis can be conducted by taking the linearization error into consideration.
Constraint satisfiability and Forward Invariance: In the pick-and-place task, an MPC task controller is tracked by a safe controller. The safe controller is realized with a safety index (also known as the barrier function), which guarantees collision avoidance with mathematical proofs [77, 78]: it ensures the system state always satisfies $\min\{d_{\text{min}}-d(x),100\cdot(0.02^{2}-d(x)^{2})-10\cdot\dot{d}(x)\}<0$, where $d_{\text{min}}$ is the allowable minimum distance between the robot and the obstacle, $d(x)$ and $\dot{d}(x)$ are the relative distance and relative velocity from the robot to the obstacle, respectively.
4.6 Transfer to real robot and different embodiments
The control system for opening doors is synthesized in simulation and is executed in both simulation and the real world. As shown in fig. 3 and fig. 4, the behavior is consistent and no sim-to-real gap is observed because the synthesized controller is model-based, closed-loop, highly explainable, and math-guaranteed. The controller can also generalize to different embodiments with the same low-level API (e.g. 7 DoF joint torque) given the model of the new embodiments, shown in fig. 7.
Balance Open Door Safe Pick&Place
Pole Mass
0.01$\sim$0.5 kg
10/10
Handle Height
0.3$\sim$0.75 m
10/10
Obstacle Position
0.01$\sim$0.3 m (y-axis)
10/10
Cart Mass
0.05$\sim$0.5 kg
10/10
Handle Radius
0.3$\sim$0.7 m
10/10
Obstacle Size
0.1$\sim$0.45 m (height)
10/10
Initial Angle
-0.5$\sim$0.5 rad
10/10
Door Mass
1$\sim$30 kg
10/10
Place Position
-0.3$\sim$0.3 m (x-axis)
10/10
Table 3: The synthesized controllers easily generalize to scenarios with different object states/attributes. The left column lists the range of the parameters for each scenario. The right column indicates the success rate (out of 10 trials) for each set of parameters.
4.7 Failure analysis
We analyze the failure cases in table 2 and summarize their reasons. 1. Mathematical error: Although LLM can give a mathematical description of the approximate dynamic model for an unseen system, it can make mistakes in math. For example, for the dynamic model synthesized for the cart-pole system, it can miss a term in the $A$ matrix, or mess up signs (use $+$ when -- is desired). 2. Failure to follow instructions: We require the LLM to provide a structured response so that a program can extract the code and plug it into the robotic system. However, sometimes LLM fails to follow the instructions, leading to responses in the wrong format. 3. Incorrect reasoning: In the open-door task, the LLM infers the target location of the door knob. However, the LLM may infer a wrong target given the environmental information. Although these are still challenging for LLMs, we believe that they can be overcome with the rapid development of LLMs in the near future.
5 Limitation and Discussion
In this work, we propose Meta-Control, a novel framework for zero-shot model-based control system synthesis using LLM, tailored for heterogeneous robotic tasks. Through both simulations and real-world tests, we demonstrate Meta-Control’s potential to expand the autonomous capabilities of robots. Despite compelling results, Meta-Control has several limitations. Such as dependency on accurate system state estimation, which may not be available in some open-world tasks, and dependency on predefined dynamic models and control templates, which may restrict adaptability to completely new heterogeneous tasks. The synthesis process also demands substantial computational resources, which could hinder real-time synthesis. Future directions include enabling automatic perception selection, incorporating learning-based templates, and accelerating synthesis speed.
6 Acknowledgement
This work was supported by the National Science Foundation under Grant No. 2144489.
Appendix A Joint torques during openning a door
As shown in fig. 8, openning a door with position control can lead to very large joint torques, leading to failure or dangerous behaviors. But with a compliant controller, the joint torques are much smaller and the door is opened successfully without damage.
Figure 8: Joint torque range during opening door for the baseline (position control) and the Meta-Control synthesized controller (stiffness control). The baseline has a huge torque because the planned trajectory is inaccurate, which leads to damage.
Appendix B Task descriptions
Open the door Opening a door is a challenge for robots because a door has a fixed swing path that must be followed exactly. As shown in fig. 3, position control can easily lead to door damage or failure of action. Therefore, it is preferable to open a door with a compliant controller. With multiple rounds of experiments, we found that Meta-Control synthesized control system usually involves a CartesianStiffnessController acting as the task controller or the tracking controller. Although the trajectory may not be perfectly aligned with the swing path, with the stiffness controller, the robot can still open the door because it complies with the force given by the door.
Wipe the board Wiping a board requires a certain amount of force to be applied on the board while moving the eraser, which involves two different objectives: position tracking and force tracking. As shown in fig. 3, the synthesized controller successfully removes the marks, while control frameworks that only consider spatial relationships are not suitable for this task because of the lack of force constraints. In most trials, Meta-Control chooses a CartesianInterpolationController as task controller, and a PoseForceController as tracking controller. The Cartesian interpolation controller plans the trajectory of the eraser, while the hybrid position/force controller tracks the trajectory while maintaining a desired force on the board to erase.
Balance the cart pole Cart pole is a classic control task that has been extensively studied. Attempts were made to synthesize a simple PID controller with LLM to balance a pole with predefined APIs where the cart can be controlled directly [1]. However, in this experiment, we use a robot arm to hold the cart and ask LLM to balance it by controlling the robot arm. This is a significantly more challenging task because only low-level APIs of the robot arm are given, and the pole is attached to the cart with a non-actuated free joint. The LLM has to understand the relationship from the arm to the cart, and from the cart to the pole. In most cases, our method chooses an LQRController as the task controller and the PoseForceController as the tracking controller. The LQR controller gives the force to be applied on the cart along the pole joint direction (y-axis) to balance the pole, and the hybrid position/force controller tracks the desired force on the y-axis while maintaining a neutral pose on the x-axis and the z-axis. Profile of the pole’s angle is shown in fig. 5, which shows that the synthesized controller efficiently balanced the pole.
The synthesized control system is described below:
$y$ $=[\text{Pole}_{\theta},\text{Pole}_{\omega},\text{Cart}_{y},\text{Cart}_{\dot{y}}]$
[…]
$x$ $=\text{Joint states},\text{EE}^{\text{target}}_{\text{force}},\text{EE}^{\text{target}}_{\text{pose}}$
[…]
where
$h(z,v)$ $=\begin{pmatrix}0&1&0&0\\ 0&0&\frac{m_{\text{pole}}g}{m_{\text{cart}}}&0\\ 0&0&0&1\\ 0&0&\frac{g(m_{\text{cart}}+m_{\text{pole}})}{l_{\text{pole}}m_{\text{cart}}}&0\end{pmatrix}z+\begin{pmatrix}0\\ \frac{1}{m_{\text{cart}}}\\ 0\\ -\frac{1}{l_{\text{pole}}m_{\text{cart}}}\end{pmatrix}v.$
Collision-free pick and place Pick and place is a very common skill in daily life. In this task, we require the robot arm to reach a goal position while maintaining whole-body collision-free. The goal position can be either the location of the object or the target location. In most cases, the LLM chooses a KinematicTrajectoryModelPredictiveController as the task controller to generate collision-free way-points for reaching the goal, and a SafeController as the tracking controller to guarantee collision-free in continuous time. In most cases, the LLM chooses a KinematicTrajectoryModelPredictiveController as the task controller to generate collision-free way-points for reaching the goal, and a SafeController as the tracking controller to guarantee collision-free in continuous time.
Appendix C Full conversation
The full conversation on skill synthesis is attached below. Some long numerical arrays are omitted for the sake of clarity. Controller templates, dynamical model templates, and input port samples are attached after the conversation.
C.1 Design via Templates
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/design.md
C.2 Interface Alignment
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/interface_alignment.md
C.3 Parameter Optimization
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/parameter_optimization.md
C.4 Model template summary
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]yamlprompt/dynamic_model_summary.yaml
C.5 Controller template summary
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]yamlprompt/controller_summary.yaml
C.6 Skill input summary
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]yamlprompt/skill_input_summary.yaml
C.7 Input port sample
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]yamlprompt/input_port_sample.yaml