### v3

# Tianhao Wei∗,1, Liqian Ma∗,1,2, Rui Chen1, Weiye Zhao1, Changliu Liu1

Abstract
The requirements for real-world manipulation tasks are diverse and often conflicting; some tasks require precise motion while others require force compliance; some tasks require avoidance of certain regions, while others require convergence to certain states. Satisfying these varied requirements with a fixed state-action representation and control strategy is challenging, impeding the development of a universal robotic foundation model. In this work, we propose Meta-Control, the first LLM-enabled automatic control synthesis approach that creates customized state representations and control strategies tailored to specific tasks. Our core insight is that a meta-control system can be built to automate the thought process that human experts use to design control systems. Specifically, human experts heavily use a model-based, hierarchical (from abstract to concrete) thought model, then compose various dynamic models and controllers together to form a control system. Meta-Control mimics the thought model and harnesses LLM’s extensive control knowledge with Socrates’ “art of midwifery” to automate the thought process. Meta-Control stands out for its fully model-based nature, allowing rigorous analysis, generalizability, robustness, efficient parameter tuning, and reliable real-time execution.
Keywords: embodied agent, model-based control, LLM, manipulation
Figure 1: Opening the door, moving safely, and balancing are heterogeneous skills and have conflicting requirements that can be difficult to satisfy with a fixed control strategy and fixed state-action representation. Meta-Control addresses the challenge with a composable hierarchical control formulation and LLM, enabling automatic model-based control synthesis. The synthesized skill has customized state action representations, dynamic models, and controllers that perfectly align with the objectives and requirements.
1 Introduction
From control (for trajectory tracking) to learning (for trajectory generation), robotic systems have gained higher autonomy over the years. However, human designers still play an indispensable role in the creation and tuning of these control and learning algorithms in a case-by-case manner, which prevents the realization of general-purpose robots. For robots to better serve and assist human users in a variety of tasks, there is a pressing need for higher autonomy that allows robots to autonomously synthesize skills $S$ based on language instructions $\ell$ to any given task, without manual tuning. Some recent attempts have been made towards this goal using LLM. Such as automatically selecting a skill $S_{i}$ from a predefined skill library $\{S_{1},S_{2},\cdots\}$ [1, 2, 3], leveraging generic parameterized policies $S_{\theta}$, where $\theta:=LLM(\ell)$ [4, 5, 6], and reinforcement learning (RL) with LLM generated rewards: $S:=RL(LLM(\ell))$ [7, 8, 9]. However, the journey towards general-purpose robots, capable of adapting to arbitrary tasks without manual tuning, remains incomplete.
A major challenge for general-purpose robots is that real-world tasks have inherently diverse and often conflicting requirements that are difficult to satisfy with a fixed control strategy. For example, as illustrated in fig. 1, moving safely requires precise movements for collision avoidance, necessitating safe position control. Conversely, position control is unsuitable for opening the door due to the difficulty of planning a trajectory that perfectly aligns with the door’s swing path, making force compliance control more favorable. Similarly, balancing demands controllers that ensure the pole’s convergence (a position-attracting goal), which opposes collision avoidance (a position-avoiding goal). Although visuomotor methods can potentially address heterogeneous tasks, they often lack explainability and reliability. Predefined action primitives may cover different task types but require manual construction and have limited applicability and scalability. These challenges restrict existing methods from generalizing to various open-world manipulation tasks with varying constraints.
To advance towards a general-purpose robot, we propose Meta-Control, a novel framework $F$ that automatically synthesizes model-based robot skills from instructions for diverse open-world tasks: $S:=F(\ell)$. Our core insight is that higher level autonomy can be achieved by creating a meta-control system to automate the thought process that human experts use to design control systems, with the help of LLMs. More specifically, human experts design different representations and control strategies for different tasks following a model-based, hierarchical (e.g., from abstract to concrete) thought model, and then compose various dynamic models and controllers into an integrated control system. By mimicking the thought model and the thought process, the autonomous system, Meta-Control, can gain greater coverage in robotics problems with good performance.
Based on these insights, we formulate robot skill design as a hierarchical control system synthesis problem and leverage LLM to customize representations and control strategies. The hierarchy involves a high-level task space and a low-level tracking space. The task space is an abstracted, intuitive space for accomplishing the task (e.g., Cartesian space or gripper pose space) while the tracking space usually represents the robot state space where low-level constraints can be specified and task-level commands are followed. Meta-Control first let the LLM determine the task and tracking spaces, then design the corresponding dynamic models and controllers utilizing predefined but abstract templates. The abstract model/controller template grounds the behavior of the synthesized system while maintaining high flexibility. We design a set of guidance and checklist of error in the prompt to help the LLM avoid potential errors and produce the correct code, just as how Socrate helps people draw their own conclusions.
Meta-Control offers several benefits: 1. It enables the synthesis of challenging heterogeneous robotic skills for unseen tasks, allowing each task to be accomplished with the most suitable representation and control strategy tailored to task-specific requirements. 2. Unlike previous work that primarily utilizes spatial priors from LLMs (e.g., object localization, affordance), Meta-Control leverages the internalized control knowledge of LLMs which encompasses decades of modeling and optimization efforts for various tasks and skills. 3. The synthesized control system is fully model-based which brings robustness, generalizability, efficient parameter tuning, and formal guarantees (e.g., safety and stability), leading to reliable and trustworthy execution.
2 Related work
We review four closely related approaches to achieve higher autonomy: skill libraries, optimization, end-to-end methods, and model-based control synthesis. Meta-Control integrates aspects of these approaches, offering a flexible, robust, and explainable solution for diverse real-world tasks.
Skill libraries with LLM methods utilize predefined libraries of motion primitives, enabling diverse control strategies through high-level APIs. LLMs can dynamically combine these libraries for task execution [1, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Skill libraries are often constructed using behavior cloning, reinforcement learning, or bootstrapping [5, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. However, these motion primitives are manually constructed and task-specific. In contrast, Meta-Control is capable of synthesizing new heterogeneous skills on the fly.
Optimization-based methods use a fixed optimization framework such as MPC or RL with mutating hyper-parameters, including objectives [30, 31, 32, 33, 7, 8], constraints [34, 4, 35], or dynamic models [36, 37, 38, 39, 40, 41]. These hyper-parameters can be inferred by LLMs [4, 33, 7, 8] or learned from data [36, 30, 31]. Hierarchical optimization often use LLM inferred high-level commands for an instruction-following controller [42, 29]. A key limitation is that the chosen framework constrains the method’s capability. An MPC generating end-effector actions cannot produce force-compliant control in joint space, and RL-based methods are unsuitable for online skill synthesis. Meta-Control, however, dynamically generates flexible policies.
End-to-end models directly map perceptions and instructions to robot actions, using VLMs [6, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] or energy functions and affordance maps [53, 54, 55, 56]. Diffusion models have also been used to learn from demonstration [57, 58, 20], enabling multimodal action distribution. End-to-end methods often output Cartesian actions, requiring additional motion planning for joint-level movement, and they are data-hungry. Furthermore, these methods lack robustness and explainability. Meta-Control is data-efficient, integrating LLM with model-based control strategies to ensure flexibility, robustness, and explainability.
Model-based control synthesis designs explainable systems with rigorous guarantees [59, 60, 61, 62, 63, 64, 65]. Hierarchical formulations simplify control synthesis for complex systems [62, 66, 67, 68, 69, 70, 71]. However, model-based synthesis is often task-specific and lacks generalizability. Meta-Control overcomes this limitation with the help of LLM.
Figure 2: Comparison between Meta-Control and a trajectory planning-based method on a real robot for wiping the board and opening the cabinet. The trajectory-based method fails to erase the mark because it neglects force requirements. Opening the cabinet with a trajectory-based controller leads to cabinet displacement because the planned trajectory does not precisely align with the door’s swing path, which may damage the door if the cabinet is fixed. In contrast, Meta-Control addressed these challenges with properly customized control systems.
3 Method
Figure 3: Overview of Meta-Control: The user only needs to provide a skill description. Meta-Control then leverages the control knowledge of LLMs to synthesize skills through a three-level pipeline: strategy level, data flow level, and parameter level. For each level, we have designed a generic prompt with placeholders, which are dynamically replaced with user input or code extracted from the LLM response during runtime, utilizing a code extractor to make the prompt task-specific. The extracted code is also used to construct the control system. At each level, if the LLM-generated code results in an error, a reflection phase is initiated. We have embedded design principles and checklists of common errors within the design and reflection prompts to assist the LLM in producing correct code. The generic prompt design with placeholders allows Meta-Control to generalize to unseen tasks without modification.
In this work, we focus on synthesizing robot skills, defined as unit actions (e.g., grab the eraser, erase the marks) from robot tasks instructed via free-form language $\mathcal{L}$ (e.g., clean the whiteboard) [4]. We assume that the decomposition from task to skills: $\mathcal{L}\to{\ell_{1},\ell_{2},\dots,\ell_{n}}$ is given by a task-level planner, which can be LLM-based or search-based [72]. Our focus is to synthesize a control system to perform a skill described by $\ell$.
Designing control systems from scratch for open-world skills is very difficult even for humans because the design space is infinite. Mimicking the thought model of control experts, we propose a composable hierarchical control formulation and introduce the Meta-Control pipeline.
3.1 Meta-Control Formulation: Composable Hierarchical Control
Hierarchical Control The bilevel hierarchical control formulation can represent various skills through a high-level controller in an intuitive task space and a low-level controller that tracks the high-level control in the robot state space. Formally, we define
[…]
where $x$ is the system state, $u$ is the state space control input, $y$ is the output or measurement of the system, $z$ is the state of the task space, and $v$ is the task space control input. The task space contains intuitive and high-level states, such as the gripper poses for robot arms and the center-of-mass for quadrupeds. $x$ and $z$ may not be directly observable, but $y$ has to be measurable or extractable from perception. $f,h$ are dynamic models of the corresponding system. We assume $g$, the perception algorithms, is given in this work, but $g$ could also be designed. The task space can sometimes be omitted when it coincides with the state space, such as when a robot arm skill directly specifies joint goals. We denote the task space controller by $\pi_{v}(y)$, and the tracking space controller by $\pi_{u}(y,v)$.
Our goal is to use LLM to 1) infer the objectives and constraints of the task; 2) design the proper task state $z$; 3) design the dynamic models $h(z,v)$ and $f(x,u)$; 4) design the task and tracking controller $\pi_{v}(y)$, $\pi_{u}(y,v)$ to achieve the desired performance.
Composable Design Designing $z$, $h$, $f$, $\pi_{v}$, and $\pi_{u}$ directly is still very challenging due to infinite possible spaces and dynamic models and the need for an accurate and deep understanding of the robotic system. Similar to humans, LLMs excel in making intuitive decisions, but are relatively poor at extensive reasoning, large-scale design, and implementation [73]. Furthermore, LLM generated code cannot guarantee constraint satisfaction. Therefore, we introduce model and controller templates, which enable composable design and guaranteed constraint satisfaction.
Templates are predefined object classes that need to be instantiated with concrete arguments. For example, we offer a dynamic model template called LinearModel which can be instantiated by passing in four matrices $A,B,C,D$ describing a linear system, and a controller template LQRController requires matrices $Q,R$ and vectors $x_{0},u_{0}$. Templates differ from motion primitives in that they are abstract and generic. The usage of templates grounds the behavior of the system, greatly increasing explanability while maintaining flexibility.
3.2 Meta-Control Pipeline: Generic Prompts with Multi-Level Synthesis
As illustrated in fig. 3, we designed a 3-level pipeline for Meta-Control (Meta-Control) to generate hierarchical control systems using generic prompts that adapt to unseen tasks without modification. The user provides a task description, and Meta-Control interacts with an LLM in a Socratic manner to synthesize the control system. A detailed query and response can be found in appendix C.
Art of Midwifery LLMs often struggle to write fully functional code without errors. To guide it, we implemented a method similar to Socratic questioning, where instead of directly giving solutions, we prompt the LLM to “give birth” to correct code through structured guidance. Meta-Control interacts with the LLM through three dialogues, each representing a level of the pipeline, and after each, it generates a summary that informs the next step. These dialogues follow a structured template, with task-specific placeholders, filled in by user input or code extracted from the LLM’s responses using a code extractor. This process ensures the pipeline can adapt to unseen tasks, simply by providing task descriptions. Each dialogue follows specific design principles, such as defining control spaces based on available measurements, designing the task controller with objectives in mind, and aligning the tracking controller with task requirements. If the generated code encounters errors, a reflection phase, depicted in fig. 3., is triggered. This phase uses a checklist of common LLM errors and related error messages to correct them. For example, Drake’s [74] unusual rotation-translation order is a common source of errors. A detailed breakdown of prompt design techniques is provided in appendix B.
Strategy Level: Design via Template Meta-Control first queries the LLM to 1) identify the appropriate task space, objectives, and constraints; 2) selects the models and controllers from the template library, and 3) instantiate them with the correct parameters based on the task’s requirements. The process follows a logical sequence: defining the task controller, selecting the task model, and then designing the tracking model and controller. The LLM’s responses follow a strict format, allowing us to extract the generated code and evaluate it.
Data Flow Level: Composition via Semantic Understanding emplate inputs and outputs often differ in format and semantic meaning, so the LLM must align these interfaces to ensure unimpeded data flow. Meta-Control queries the LLM to implement two objects: 1) a task input converter that composes the task controller’s input from available measurements $y$; 2) a tracking input converter that composes the tracking controller’s input from $y$ and the task control $v$. F In the balance cart-pole example, the task input converter extracts the 4D input for the LQRController from 20+ available measurements. It then passes the task control (a scalar force) to the tracking input converter, which prepares the inputs for the HybridPositionForceController. The tracking input converter has to pad the task control and combine it with available measurements to form a 12D input. A checklist of common errors helps correct interface mismatches, such as dimensionality or data type issues.
Parameter Level: Alignment via Few-Shot Optimization The designed system may degenerate when deployed because the inferred parameters may not align the reality. To address this, Meta-Control enables few-shot optimization of system parameters by querying the LLM on 1) key parameters affecting behavior, 2) performance metrics, and 3) strategies for adjusting parameters based on past trajectories. For instance, in the cart-pole task, the LLM identifies the LQR controller’s $Q$ and $R$ parameters as crucial, using the cart’s position and pole’s angle as metrics. The process is efficient because LLMs have internalized parameter-performance relationships in common dynamic systems. A list of possible root causes (e.g., initialization errors or suboptimal parameters) of the degenerated performance helps the LLM diagnose problems more effectively.
4 Experiment
Figure 4: Five manipulation tasks that have inherently different challenges and requirements. For example, the balance task requires an accurate and high-frequency feedback controller. The safe pick and place task requires guaranteeing collision avoidance for the whole robot arm. The open door task requires properly handling articulated objects; and the executed trajectory has to perfectly match the swing path.
Figure 5: Meta-Control can automatically identify hyper-parameters that require tuning and tune them to accomplish challenging tasks. The figure shows the trajectory of the arm-held cart-pole system before and after tuning the synthesized controller. The hyper-parameters $Q=$diag$(10,1,100,1)$, $R=0.01$ are chosen and tuned by the LLM with only 2 rounds of trial-and-error.
The experiment is designed to manifest the following features. We first show that Meta-Control enables synthesizing skills for heterogeneous requirements by exploiting the control knowledge of LLM. Then we did an ablation study to show that all three levels of the Meta-Control pipeline improve the success rate; Finally, we reveal the benefit of the model-based design, including robustness, generalizability, and formal guarantees.
Most of the experiments are conducted on four representative tasks. These tasks cover constraints that appear frequently in the real world, such as constraints on position, velocity, force, compliance, convergence, and safety. We implement our pipeline with Drake[74], a framework designed for model-based control. For the hardware experiment, we used a Kinova Gen3 robot arm. The language model is GPT 4.0 with a default temperature $1.0$.
4.1 Meta control enables synthesizing challenging skills
As shown in fig. 2 and fig. 4, Meta-Control successfully synthesizes controllers for various challenging tasks with inherently different requirements, both in simulation and in the real world. Figure 5 shows that Meta-Control can identify core metrics and critical parameters that affect performance, then efficiently and effectively tune the parameters based on the metric through trial and error. A detailed description of task challenges and synthesized control systems can be found in section A.1. The complete skill synthesis process can be found in appendix C.
Meta-Control exploits dynamics priors In contrast to previous work that focuses more on spatial relationship priors [1, 4, 5], Meta-Control exploits the LLM’s internalized knowledge of dynamics, enabling rigorous synthesis of high-performance controllers. In the balance cart pole task, LLM designs the task space dynamics $h(z,v)$ with a linear approximation in the form of $\dot{z}=Az+Bv$ around the upright position of the pole, where $A$ and $B$ are correctly given by the LLM: $A=[[0,1,0,0];[0,0,\frac{m_{\text{pole}}g}{m_{\text{cart}}},0];[0,0,0,1];[0,0,\frac{g(m_{\text{cart}}+m_{\text{pole}})}{l_{\text{pole}}m_{\text{cart}}},0]]$, $B=[0;\frac{1}{m_{\text{cart}}};0;-\frac{1}{l_{\text{pole}}m_{\text{cart}}}]$, where $m$ denotes mass, $l$ denotes length, $g$ is the acceleration of gravity. Detailed description of the synthesized control system can be found at section A.1.
Method API API + Hierarchy API + Template Meta-Control
Balance design 30% 90% 60% 100%
implementation 0% 30% 20% 90%
execution 0% 0% 0% 70%
Open door design 40% 50% 60% 100%
implementation 10% 20% 10% 100%
execution 0% 0% 0% 80%
Safe Pick&place design 0% 0% 40% 90%
implementation 0% 0% 0% 90%
execution 0% 0% 0% 90%
Table 1: The ablation study shows that hierarchical formulation (Hierarchy) and templates-based synthesis (Templates) both improves the success rate, compared to directly synthesis with only the low-level API. API + Template can be viewed as a generalized motion-primitive method.
4.2 Ablation study of the Meta Control pipeline
To demonstrate the necessity of the hierarchy formulation and simplification using templates, we show the success rate of control system synthesis with different ways to query the LLM. Meta-Control use the query: “output is 7-DoF torque, input is …” (API) + “consider a task space and a tracking space…” (Hierarchy) + “use available templates” (Template). We test the success rate in 3 stages: design (via templates), composition (interface alignment), and execution (after few-shot optimization). We say that a design is successful if the LLM-designed control system has the potential to perform the skill judged by an expert. We say that the implementation is successful if the system can run without errors. We say that the execution is successful if the control system finishes the skill as desired. We repeat each task $10$ times to compute the success rate. Randomness is caused by the LLM. As shown in table 1, we can see that the success rates of all steps in the baseline are lower than Meta-Control. Although LLM gives reasonable architectures to finish the task in the Balance and Open Door task, the LLM fails to provide correct code and parameters to realize the control system due to the complexity of the system and the huge design space. With all the modules, we achieve the highest success rate for all tasks. The Safe Pick&Place task is especially difficult for the baseline because the baseline methods, even though prompted to avoid collision, were unable to successfully design a controller that can avoid collision continuously.
4.3 Meta-Control brings benefits of model-based design
The model-based nature of the synthesized controller brings a variety of benefits, such as robustness, explainability, generalizability, and rigorous analysis.
Generalization to different attributes/states Meta-Control synthesized control systems can easily generalize to scenarios of different attributes/states due to the model-based nature. Given a successfully synthesized control system, we test different attributes/states and calculate the success rate. The range of change and the results are shown in table 2. Examples are shown in fig. 7. Meta-Control achieved a 100% success rate for all scenarios.
Balance Open Door Safe Pick&Place
Pole Mass
0.01$\sim$0.5 kg
10/10
Handle Height
0.3$\sim$0.75 m
10/10
Obstacle Position
0.01$\sim$0.3 m (y-axis)
10/10
Cart Mass
0.05$\sim$0.5 kg
10/10
Handle Radius
0.3$\sim$0.7 m
10/10
Obstacle Size
0.1$\sim$0.45 m (height)
10/10
Initial Angle
-0.5$\sim$0.5 rad
10/10
Door Mass
1$\sim$30 kg
10/10
Place Position
-0.3$\sim$0.3 m (x-axis)
10/10
Table 2: The synthesized controllers easily generalize to scenarios with different object states/attributes. The left column lists the range of the parameters for each scenario. The right column indicates the success count of 10 trials for each set of parameters.
Rigorous Formal Analysis The model-based design allows rigorous formal analysis for a variety of properties, making autonomous certification by LLM possible. For example, we can guarantee the convergence for the cart-pole stask by solving the Riccati equation, and ensure collision avoidance for the pick-and-place task by certifying a safety index. Details can be found in section A.3
Transfer to real robot and different embodiments The control system for opening door is synthesized in simulation and is executed both in simulation and in the real world. As shown in fig. 2 and fig. 4, the behavior is consistent and no sim-to-real gap is observed because the synthesized controller is model-based, closed-loop, highly explainable, and math-certified. The controller can also generalize to different embodiments with the same low-level API (e.g. 7 DoF joint torque) given the model of the new embodiments as shown in fig. 7.
Figure 6: Meta-Control synthesized control systems are robust to attribute/state changes because of the model-based design.
Figure 7: Meta-Control generalizes to different embodiments because the synthesis is fully model-based. Furthermore, a controller synthesized on Kinova can transfer to Franka Pranda simply by replacing the robot dynamic model thanks to the model-based nature.
4.4 Failure analysis
We analyze the failure cases in table 1 and summarize their reasons. 1. Mathematical error: Although LLM can give a mathematical description of the approximate dynamic model for an unseen system, it can make mistakes in math. For example, for the dynamic model synthesized for the cart-pole system, it can miss a term in the $A$ matrix, or mess up signs (use $+$ when -- is desired). 2. Failure to follow instructions: We require the LLM to provide a structured response so that a program can extract the code and plug it into the robotic system. However, sometimes LLM fails to follow the instructions, leading to responses in the wrong format. 3. Incorrect reasoning: In the open-door task, the LLM infers the target location of the door knob. However, the LLM may infer a wrong target given the environmental information. Although these are still challenging for LLMs, we believe that they can be overcome with the rapid development of LLMs in the near future.
5 Limitation and Discussion
In this work, we present Meta-Control, a novel framework for automatic model-based control system synthesis using LLMs, designed for heterogeneous robotic tasks. Simulations and real-world tests demonstrate its potential to enhance robotic autonomy. However, Meta-Control has limitations, such as reliance on accurate system state estimation and predefined models, which limit adaptability to new tasks. The LLM can also make suboptimal decisions, like selecting PID over LQR, complicating parameter tuning. Future improvements include better reflection across the pipeline, stronger reasoning capabilities, faster synthesis, and the integration of learning-based templates and automatic perception selection.
6 Acknowledgement
This work was supported by the National Science Foundation under Grant No. 2144489.
Appendix A Experiment details
A.1 Task descriptions
Task Challenge Meta-Control designed system
Open the door The robot trajectory must perfectly align with the door’s swing path. position control can easily lead to damage or failure. CartesianTrajectoryController + CartesianStiffnessController, allowing imperfect trajectory planning and tracking with compliant behavior to avoid damage or failure.
Balance the cart pole The pole is non-actuated. The system is sensitive, requiring high-frequency feedback and convergence guarantee. LQRController + HybridPositionForceController. The LQR controller gives the force to be applied on the cart, and the hybrid position/force controller tracks the desired force on the y-axis while maintaining a neutral pose on the x-axis and the z-axis.
Collision-free pick and place whole-body collision free in continuous time during the whole task. KinematicTrajectoryMPC + SafeController, allowing discrete-time planning and continuous-time whole-body collision-free tracking.
Wipe the whiteboard Two different objectives: tracking position and maintaining force CartesianTrajectoryController + HybridPositionForceController, allowing position tracking while maintaining a desired force on the whiteboard.
Table 3: Experiment tasks, challenges and Meta-Control synthesized controllers..
Open the door Opening a door is a challenge for robots because a door has a fixed swing path that must be followed exactly. As shown in fig. 2, position control can easily lead to door damage or failure of action. Therefore, it is preferable to open a door with a compliant controller. With multiple rounds of experiments, we found that Meta-Control synthesized control system usually involves a CartesianStiffnessController acting as the task controller or the tracking controller. Although the trajectory may not be perfectly aligned with the swing path, with the stiffness controller, the robot can still open the door because it complies with the force given by the door.
Wipe the board Wiping a board requires a certain amount of force to be applied on the board while moving the eraser, which involves two different objectives: position tracking and force tracking. As shown in fig. 2, the synthesized controller successfully removes the marks, while control frameworks that only consider spatial relationships are not suitable for this task because of the lack of force constraints. In most trials, Meta-Control chooses a CartesianInterpolationController as task controller, and a PoseForceController as tracking controller. The Cartesian interpolation controller plans the trajectory of the eraser, while the hybrid position/force controller tracks the trajectory while maintaining a desired force on the board to erase.
Balance the cart pole Cart pole is a classic control task that has been extensively studied. Attempts were made to synthesize a simple PID controller with LLM to balance a pole with predefined APIs where the cart can be controlled directly [1]. However, in this experiment, we use a robot arm to hold the cart and ask LLM to balance it by controlling the robot arm. This is a significantly more challenging task because only low-level APIs of the robot arm are given, and the pole is attached to the cart with a non-actuated free joint. The LLM has to understand the relationship from the arm to the cart, and from the cart to the pole. In most cases, our method chooses an LQRController as the task controller and the PoseForceController as the tracking controller. The LQR controller gives the force to be applied on the cart along the pole joint direction (y-axis) to balance the pole, and the hybrid position/force controller tracks the desired force on the y-axis while maintaining a neutral pose on the x-axis and the z-axis. Profile of the pole’s angle is shown in fig. 5, which shows that the synthesized controller efficiently balanced the pole.
The synthesized control system is described below:
$y$ $=[\text{Pole}_{\theta},\text{Pole}_{\omega},\text{Cart}_{y},\text{Cart}_{\dot{y}}]$
[…]
$x$ $=\text{Joint states},\text{EE}^{\text{target}}_{\text{force}},\text{EE}^{\text{target}}_{\text{pose}}$
[…]
where
$h(z,v)$ $=\begin{pmatrix}0&1&0&0\\ 0&0&\frac{m_{\text{pole}}g}{m_{\text{cart}}}&0\\ 0&0&0&1\\ 0&0&\frac{g(m_{\text{cart}}+m_{\text{pole}})}{l_{\text{pole}}m_{\text{cart}}}&0\end{pmatrix}z+\begin{pmatrix}0\\ \frac{1}{m_{\text{cart}}}\\ 0\\ -\frac{1}{l_{\text{pole}}m_{\text{cart}}}\end{pmatrix}v.$
Collision-free pick and place Pick and place is a very common skill in daily life. In this task, we require the robot arm to reach a goal position while maintaining whole-body collision-free. The goal position can be either the location of the object or the target location. In most cases, the LLM chooses a KinematicTrajectoryModelPredictiveController as the task controller to generate collision-free way-points for reaching the goal, and a SafeController as the tracking controller to guarantee collision-free in continuous time. In most cases, the LLM chooses a KinematicTrajectoryModelPredictiveController as the task controller to generate collision-free way-points for reaching the goal, and a SafeController as the tracking controller to guarantee collision-free in continuous time.
A.2 Joint torques during openning a door
As shown in fig. 8, openning a door with position control can lead to very large joint torques, leading to failure or dangerous behaviors. But with a compliant controller, the joint torques are much smaller and the door is opened successfully without damage.
Figure 8: Joint torque range during opening door for the baseline (position control) and the Meta-Control synthesized controller (stiffness control). The baseline has a huge torque because the planned trajectory is inaccurate, which leads to damage.
A.3 Formal Analysis of Properties
Convergence and Stability: For the balance cart-pole task, the convergence can be guaranteed by solving the Riccati equation for the LQR controller. The closed-loop system matrix $A-BK$ has the following four eigenvalues: $-412.29,-9.925,-1.502+1.175j,-1.502-1.175j.$ All of them have negative real parts, which means that the system is guaranteed to converge. More rigorous analysis can be conducted by taking the linearization error into account.
Constraint satisfiability and Forward Invariance: In the pick-and-place task, an MPC task controller is tracked by a safe controller. The safe controller is realized with a safety index (also known as the barrier function), which guarantees collision avoidance with mathematical proofs [75, 62]: it ensures the system state always satisfies $\min\{d_{\text{min}}-d(x),100\cdot(0.02^{2}-d(x)^{2})-10\cdot\dot{d}(x)\}<0$, where $d_{\text{min}}$ is the allowable minimum distance between the robot and the obstacle, $d(x)$ and $\dot{d}(x)$ are the relative distance and relative velocity from the robot to the obstacle, respectively.
A.4 Runtime and parameter optimization rounds
In our experiments, it usually takes GPT-4o 1-3 minutes to synthesize a skill, depending on the capacity of the servery.
Appendix B Prompt Design
We use a variety of techniques to improve the quality of the full dialogue. First, we introduce some techniques in the overall design, followed by the techniques used in each step of the three-level pipeline.
B.1 Overall Design Techniques
  1. 1.
    Divide skill synthesis into three dialogues, corresponding to the three-level pipeline. This approach avoids overly long contexts, enables step-by-step reasoning, and allows reflections to focus on the current step.
  2. 2.
    At the end of each dialogue, Meta-Control asks the LLM to generate a summary to warm-start the next dialogue, including the decisions made in the current step and all necessary information for the next step.
  3. 3.
    Use a generic prompt with placeholder design to make the prompt zero-shot generalizable while remaining task-specific. Each placeholder is uniquely identified by a <step_name>. The placeholders are dynamically replaced by the LLM generated code.
  4. 4.
    Provide guidance and examples to help the LLM generate code following a specific format. The code should start with a <step_name>, e.g., <task_model>, followed by a Python code block encompassed by triple backticks (‘‘‘).
  5. 5.
    Use <step_name> to insert LLM-generated code into the control system. Our code extractor uses regular expressions to search for <step_name>, extract code from the LLM output, and then insert the code into the control system based on <step_name>.
  6. 6.
    Use a closed-loop design procedure. The LLM-generated code is executed by our program, which receives feedback, such as error messages or measurements during execution. Meta-Control then asks the LLM to improve the code based on the feedback.
  7. 7.
    Use code replacement instead of code completion to progressively construct the control system. Debugging and improvement are difficult with an incomplete control system code that is only executable after completing it with LLM-generated code, as feedback for intermediate steps would be unavailable. Therefore, we initially provide a blank but executable control system code, which outputs all zero joint torques. At each step of the synthesis, part of the code is replaced based on the identifier. We designed the control system to be always executable during the design process as long as the generated code is correct.
B.1.1 Dialogue 1: Design via Template Prompt
  1. 1.
    Use a role-play prompt (e.g., a professor in control and a proficient programmer) to invoke control knowledge and coding skills.
  2. 2.
    Provide some optional principles to assist the design, such as “design the task controller on simplified dynamics” and “use the task controller to accomplish the primary objective while using the tracking controller to satisfy constraints or secondary objectives.”
  3. 3.
    Use model templates and controller templates to ground the design space. The templates greatly reduce the design space and anchor the behavior of the synthesized control system, making it explainable and provable.
  4. 4.
    Provide all available measurements and their samples. The sample measurements give the LLM a precise idea of what inputs are available and their format, helping it to decide on the appropriate task controller and tracking controller.
  5. 5.
    Provide an example of the code that the LLM should return, including the <step_name> and how to initialize the chosen models and controllers.
  6. 6.
    Provide a checklist of requirements that the LLM should satisfy to reduce errors, such as “choices have been made for all steps” and “<step_name> must be included before each code block.”
  7. 7.
    Use reflections. Meta-Control extracts and evaluates the code from the LLM response, then informs the LLM if the initialization of the chosen models and controllers causes any errors, allowing the LLM to decide if any fixes or improvements are needed. This process is repeated until there are no errors or the maximum iteration is reached.
  8. 8.
    Ask the LLM to generate a summary of the design to warm-start the composition dialogue. The summary includes the chosen models and controllers, a detailed description of the inputs for the controllers, the meaning of each dimension, and any other necessary information.
B.1.2 Dialogue 2: Composition via Semantic Understanding
  1. 1.
    Use a role-play prompt as in the previous dialogue.
  2. 2.
    Use the design summary to warm-start the dialogue, providing the task, the chosen controllers and models, their input and output ports, and how they should be combined into a control system to complete the task.
  3. 3.
    Provide examples of how to access the values of the input and output ports of a model or controller.
  4. 4.
    Provide an example of the input converter, demonstrating how to connect the input and pad zeros when necessary. The input converter is a set of callback functions that assign values to the input based on the output of other modules.
  5. 5.
    Provide tips and potential errors to avoid when implementing the input converter. For example, “Ensure the dimensions of the output match the port. The output value may only be partially available from the input port,” and “For pose, velocity, and force vectors, rotation always comes first.”
  6. 6.
    Use partial updates. Because input converters involve many functions, if an error is detected during reflection, Meta-Control only updates the function that the LLM identifies as causing the error, instead of updating all functions, which may cause more errors if the code is not exactly the same.
  7. 7.
    Ask the LLM to generate a summary of the data flow to warm-start the parameter tuning, including parameters of the controller, parameters of the model, input converter structures, and potential performance metrics.
B.1.3 Dialogue 3: Parameter Optimization
  1. 1.
    Use a role-play prompt as in the previous dialogues.
  2. 2.
    Provide a list of code blocks that the LLM can update to optimize the parameters, covering models, controllers, and input converters.
  3. 3.
    Warm-start the dialogue with the data flow summary, providing a detailed implementation of the control system, a description, and samples of the module inputs.
  4. 4.
    Provide a guide for parameter optimization: identify key parameters, identify performance metrics from task state and task control, and then determine if the time series of the metrics is desired.
  5. 5.
    Provide a checklist of errors to identify the cause of degraded performance, including definition, initialization, implementation, and suboptimal parameters.
  6. 6.
    Iteratively optimize the parameters with reflections. Every time the LLM updates the code or parameters, Meta-Control evaluates the new system, provides a time series of the chosen metrics, and allows the LLM to decide if the performance is desired. This process is repeated until the performance is satisfactory.
Appendix C Full conversation
The full conversation on skill synthesis is attached below. Some long numerical arrays are omitted for the sake of clarity. The controller templates, the dynamical model templates, and the input port samples are attached after the conversation.
C.1 Balance
C.1.1 Design via Templates
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/design.md
C.1.2 Composition via Semantic Understanding
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/interface_alignment.md
C.1.3 Parameter Optimization
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/parameter_optimization.md
C.2 Pick and Place
C.2.1 Design via Templates
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/pick_and_place/pick_and_place_design.md
C.2.2 Composition via Semantic Understanding
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/pick_and_place/pick_and_place_implement.md
C.3 Open Door
C.3.1 Design via Templates
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/open_door/open_door_design.md
C.3.2 Composition via Semantic Understanding
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/open_door/open_door_implement.md
C.4 Model template summary
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]yamlprompt/dynamic_model_summary.yaml
C.5 Controller template summary
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]yamlprompt/controller_summary.yaml
C.6 Skill input summary
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]yamlprompt/skill_input_summary.yaml
C.7 Input port sample
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]yamlprompt/input_port_sample.yaml